6-Find the general solution of the following system. Write the solution in matrix form. (7 Points) X - 2x + y y = 8x - 5y
Answer #1

x = 2x + y y = Zr-5y writing it in matrix form. Let X= [3] 2X1 x = AX, where, Az ( i< (3]- [3 ][ ] 21 -5 - 2x2 Caracteristic=> X = çe 8 by] + zesti writing it in matrix form X = { }} = (-6, e-bt + C e 8c, ett + chelt 3t x =[%] = [. с 2. [ este C2 3t